# Categorical Time-Reversal Symmetries

### 30 Mar 2026 v1

@@ IV.2.3 Examples: $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$
According to Galois descend, $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$ can be understood as $\mathsf{Rep}(G)$ equipped with a ${\mathbb{Z}}_{2}^{T}$-action. Here the action is straightfoward: $T:R\mapsto R^{*}$ where $R^{*}$ is the complex conjugate of $R$. If $R$ is not invariant under $T$, meaning that $R^{*}$ is not equivalent to $R$, then the invariant combination is $R\oplus R^{*}$, which descends to a complex simple object in $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$. On the other hand if $R\simeq R^{*}$, then $R$ descends to either a real object or a quaternion object in $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$. Which case is true can be determined by the the Frobenius–Schur indicator, defined as
@@ IV.2.4 Module categories over $\mathsf{Rep}(G^{T})$
that is, the category of projective representations of $K$ with projective phase $\psi$. The indecomposable modules over $\mathsf{Rep}(G^{T})$ has essentially the same structure. An indecomposable module is labeled by a pair $(K^{T},\psi)$ where $K^{T}<G^{T}$ and $\psi$ is a twisted 2-cocyle of $K^{T}$. The corresponding module category is
@@ IV.4 The $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$ Categories and Time-Reversal SSB
We note that it is not always possible to “add time-reversal” to an internal generalized symmetry. The above construction requires the existence of a categorical ${\mathbb{Z}}_{2}^{T}$-action on the internal generalized symmetry $\mathcal{C}$, which may not always exist. For instance, the existence of an anti-linear action requires $\mathcal{C}$ to be equivalent to its complex conjugation $\mathcal{C}^{*}$. Consider for instance the fusion category $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$ with a nontrivial cocyle $\omega\in H^{3}({\mathbb{Z}}_{3},U(1))={\mathbb{Z}}_{3}$, then $(\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega})^{*}\simeq\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega^{-1}}$, which is not equivalent to $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$. This means a system with $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$ symmetry can not be further enriched with ${\mathbb{Z}}_{2}^{T}$-symmetry.
@@
Consider a (1+1)(1+1)d system/phase whose topological defects form a complex fusion category $\mathcal{C}$. The phase could be either gapless(say a CFT) or gapped(say an SPT protected by some internal generalized symmetry). Here $\mathcal{C}$ is not interpreted as a symmetry category but the category of IR symmetric defects in the phase, such as topological defects in a CFT, or topological defects in a gapped phases protected by some internal symmetry. Now assume that the system is time-reversal symmetric, then there must exist a consistent categorical ${\mathbb{Z}}_{2}^{T}$-action on the complex fusion category $\mathcal{C}$. Then the category of time-reversal symmetric defects would be made of objects in $\mathcal{C}$ that are invariant under the ${\mathbb{Z}}_{2}^{T}$-action up to isomorphism. The mathematical theory of Galois descend [72] says that taking the (homotopy) fixed points of $\mathcal{C}$ under a ${\mathbb{Z}}_{2}^{T}$-action produces an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$, whose complexification recovers $\mathcal{C}$. Now consider breaking ${\mathbb{Z}}_{2}^{T}$ spontaneously in this system(a domain wall between the two vacuua), we arrive at a new phase whose category of symmetric defects is exactly $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$. This because when time-reversal is broken, the $T\in{\mathbb{Z}}_{2}^{T}$-action creates a nontrivial IR defect in the system that fuses with defects in $\mathcal{C}$ according to the ${\mathbb{Z}}_{2}^{T}$-action on $\mathcal{C}$. The process of going from an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$ to the Morita equivalent Galois-real fusion category $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$ describes the spontaneously breaking of time-reversal.
@@ V.1 The Setup
The action of $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ on the category of boundary conditions $\mathcal{M}$ provides $\mathcal{M}$ with a module category structure: $\mathcal{C}\times\mathcal{M}\to\mathcal{M}$ that is only $\mathbb{R}$-linear: If $a$ is an anti-linear line, then a scalar on the boundary gets complex conjugated when passes the junction. This means the functor
@@
For any symmetry category $\mathcal{C}$, one can define a 2-category with
  • Objects are 2D gapped phases enriched with $\mathcal{C}$.
@@
The 2-category of 2D gapped phases enriched with $\mathcal{C}$ is $\mathsf{Mod}(\mathcal{C})$.
For instance, the 2-category of 2D gapped phases with ${\mathbb{Z}}_{2}^{T}$ symmetry is $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})\simeq 2\mathsf{Vec}_{\mathbb{R}}$ which has three simple objects: $\mathsf{Vec}_{\mathbb{R}},\mathsf{Vec}_{\mathbb{C}},\mathsf{Vec}_{\mathbb{H}}$.
@@ V.2 Order Parameters and Drinfeld Center
As we have discussed for time-reversal in Section II.3.2, a pure time-reversal symmetry does not have a string order that can detect the Haldane phase. This is not special to time-reversal: any anti-linear symmetry generator does not have a string order parameter. We will argue that given this observation, the Drinfeld center $\mathcal{Z}(\mathcal{C}_{1}\oplus\mathcal{C}_{T})$ is still the correct category of all local and string order parameters of an anti-linear symmetry $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$. Drinfeld center is a notion that can be defined for any monoidal category, hence well-defined for real fusion categories, and in particular for Galois-real fusion categories. For a categorical anti-linear symmetry $\mathcal{C}:=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$, a simple object in $\mathcal{Z}(\mathcal{C})$ is again given by a pair $(x,\gamma)$, where $x\in\mathcal{C}$ and $\gamma:x\otimes y\simeq y\otimes x,\penalty 10000\ \forall y\in\mathcal{C}$ is the half-braid. However, as shown in [15], the object $x$ is in fact strictly in $\mathcal{C}_{1}$, otherwise there is no consistent half-braid $\gamma$. This is because, assuming a half-braid $\gamma$ does exists, then it gives natural isomorphism $x\otimes\mathbf{1}\simeq\mathbf{1}\otimes x$, which identifies the left and right actions by scalars, implying $x$ must be a linear(Galois-trivial) object $x\in\mathcal{C}_{1}$. This precisely matches with the expectation that an anti-linear symmetry generator does not have a string order parameter: assume a string order parameter does exists, then it can be used to identify the action by scalars from two sides of the string order:
@@
Here we are not assuming any SymTFT interpretation of the Drinfeld center, simply viewing it as the category of 1D order parameters associated with $\mathcal{C}$. The SymTFT interpretation will be discussed in Section VI.4. However it is already clear from the discussion here that the Drinfeld center can not describe gapped phases that are not captured by string orders, such as the Haldane chain.
@@ V.3.1 Gauging between Symmetry Categories
Recall that for two complex fusion categories $\mathcal{C},\mathcal{D}$, we say they are Morita equivalent if there is some algebra $A\in\mathcal{C}$ such that $\mathcal{D}\simeq\mathsf{Bimod}_{\mathcal{C}}(A)$ are equivalent as fusion categories. In this case we say $\mathcal{D}$ is obtained by condensing $A$ in $\mathcal{C}$. Equivalently, $\mathcal{C},\mathcal{D}$ are Morita equivalent if there is a $\mathcal{C}$-module category $\mathcal{M}$ such that $\mathcal{D}\simeq\text{End}_{\mathcal{C}}(\mathcal{M})$. If $\mathcal{C},\mathcal{D}$ are symmetry categories for two internal generalized symmetries, then being Morita equivalent has the consequence that the 2-categories of gapped phases with symmetry $\mathcal{C}$ and $\mathcal{D}$: $\mathsf{Mod}(\mathcal{C}),\penalty 10000\ \mathsf{Mod}(\mathcal{D})$ are equivalent. This means we can relabel phases enriched with symmetry $\mathcal{C}$ as phases enriched with symmetry $\mathcal{D}$ in such way that does not affect any physical observables.
The Morita theory works well over $\mathbb{R}$ as well, see our discussion in Section III.3. Hence we say two anti-linear symmetries categories $\mathcal{C},\mathcal{D}$ are dual to each other if they are Morita equivalent as real fusion categories. This means there exists some algebra $A\in\mathcal{C}$ such that $\mathcal{D}\simeq\mathsf{Bimod}_{\mathcal{C}}(A)$. Notice that since both $\mathcal{C},\mathcal{D}$ are Galois-real, the algebra $A$ is necessarily in $\mathcal{C}_{1}$ according to Theorem Theorem III.4. This is the mathematical manifestation of the fact that one can not literally gauge time-reversal, or any anti-linear symmetry generator. The gauging always happens inside the linear part $\mathcal{C}_{1}$. Then if $\mathcal{C},\mathcal{D}$ are two dual anti-linear symmetries, the 2-category of gapped phases: $\mathsf{Mod}(\mathcal{C}),\penalty 10000\ \mathsf{Mod}(\mathcal{D})$ are equivalent.
@@ V.4 ${\mathbb{Z}}_{2}^{T}$ Gapped Phases Revisited
We discussed the three ${\mathbb{Z}}_{2}^{T}$ gapped phases in Section II including their module category descriptions. One thing that was not discussed there is the underlying gapped phase. Let us confirm here that the three phases has expected underlying gapped phase.
@@ V.5 General Structure of $G^{T}$ Gapped Phases
We now consider (1+1)(1+1)d gapped phases with a general group-like $G^{T}$-symmetry and with no anomaly. The symmetry category is $\mathsf{Vec}_{G^{T}}$ and $G^{T}$-gapped phases are module categories over $\mathsf{Vec}_{G^{T}}$. We have classified modules over $\mathsf{Vec}_{G^{T}}$ in Section IV.1.2. The indecomposable modules are denoted as $\mathcal{M}(K^{T},\omega)$ where $K^{T}<G^{T}$ and $\omega$ is a twisted 2-cocycle for $K^{T}$. It is clear that this matches with the standard physical classification of (1+1)(1+1)d gapped phases with $G^{T}$-symmetry. The subgroup $K^{T}$ stands for the unbroken symmetry in the phase, and $\omega$ labels the $K^{T}$-SPT phase in any of the vacuua. From the general discussion in Section V.1, the number of ground states in a phase defined by $\mathcal{M}(K^{T},\omega)$ is the number of simple objects in $\mathcal{M}(K^{T},\omega)$, which is $|G^{T}/K^{T}|$.
According to the general discussion in Sec.Section V.1, the category of symmetric defects in the phase $(K^{T},\omega)$ is given by the Morita dual real fusion category
@@
Then by Theorem Theorem III.4, this real fusion category is Galois-real exactly when $K^{T}\leq G_{0}$ is completely unitary and $\mathbb{R}$-real when $K^{T}$ contains any anti-unitary element. This also means the Morita dual categories $\mathsf{Bimod}_{\mathsf{Vec}_{G^{T}}}({\mathbb{C}}[K^{T}]^{\omega})$ for $K^{T}\leq G_{0}$ can be viewed as symmetries dual to $\mathsf{Vec}_{G^{T}}$.
@@
In the remaining of this section we study several concrete examples of gapped phases with anti-linear symmetries.
@@ V.6 ${\mathbb{Z}}_{4}^{T}$-phases
The symmetry categories $\mathsf{Vec}_{{\mathbb{Z}}_{4}^{T}}$, $\mathsf{Vec}_{{\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}^{T}}^{\alpha}$ are Mortia equivalent.
Here $\alpha$ is a nontrivial 3-cocyle given in Equation V.15. We notice this is similar to the situation with an internal ${\mathbb{Z}}_{4}$-symmetry, where it is Morita equivalent to an anomalous ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$-symmetry.

### 02 Apr 2026 v2

@@ IV.2.3 Examples: $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$
According to Galois descend, $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$ can be understood as $\mathsf{Rep}(G)$ equipped with a ${\mathbb{Z}}_{2}^{T}$-action. Here the action is straightfoward: $T:R\mapsto R^{*}$ where $R^{*}$ is the complex conjugate of $R$. If $R$ is not invariant under $T$, meaning that $R^{*}$ is not equivalent to $R$, then the invariant combination is $R\oplus R^{*}$, which descends to a complex simple object in $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$. On the other hand if $R\simeq R^{*}$, then $R$ descends to either a real object or a quaternion object in $\mathsf{Rep}(G\times{\mathbb{Z}}_{2}^{T})$. Which case is true can be determined by the Frobenius–Schur indicator, defined as
@@ IV.2.4 Module categories over $\mathsf{Rep}(G^{T})$
that is, the category of projective representations of $K$ with projective phase $\psi$. The indecomposable modules over $\mathsf{Rep}(G^{T})$ has essentially the same structure. An indecomposable module is labeled by a pair $(K^{T},\psi)$ where $K^{T}<G^{T}$ and $\psi$ is a twisted 2-cocycle of $K^{T}$. The corresponding module category is
@@ IV.4 The $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$ Categories and Time-Reversal SSB
We note that it is not always possible to “add time-reversal” to an internal generalized symmetry. The above construction requires the existence of a categorical ${\mathbb{Z}}_{2}^{T}$-action on the internal generalized symmetry $\mathcal{C}$, which may not always exist. For instance, the existence of an anti-linear action requires $\mathcal{C}$ to be equivalent to its complex conjugation $\mathcal{C}^{*}$. Consider for instance the fusion category $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$ with a nontrivial cocycle $\omega\in H^{3}({\mathbb{Z}}_{3},U(1))={\mathbb{Z}}_{3}$, then $(\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega})^{*}\simeq\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega^{-1}}$, which is not equivalent to $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$. This means a system with $\mathsf{Vec}_{{\mathbb{Z}}_{3}}^{\omega}$ symmetry can not be further enriched with ${\mathbb{Z}}_{2}^{T}$-symmetry.
@@
Consider a (1+1)(1+1)d system/phase whose topological defects form a complex fusion category $\mathcal{C}$. The phase could be either gapless(say a CFT) or gapped(say an SPT protected by some internal generalized symmetry). Here $\mathcal{C}$ is not interpreted as a symmetry category but the category of IR symmetric defects in the phase, such as topological defects in a CFT, or topological defects in a gapped phases protected by some internal symmetry. Now assume that the system is time-reversal symmetric, then there must exist a consistent categorical ${\mathbb{Z}}_{2}^{T}$-action on the complex fusion category $\mathcal{C}$. Then the category of time-reversal symmetric defects would be made of objects in $\mathcal{C}$ that are invariant under the ${\mathbb{Z}}_{2}^{T}$-action up to isomorphism. The mathematical theory of Galois descend [77] says that taking the (homotopy) fixed points of $\mathcal{C}$ under a ${\mathbb{Z}}_{2}^{T}$-action produces an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$, whose complexification recovers $\mathcal{C}$. Now consider breaking ${\mathbb{Z}}_{2}^{T}$ spontaneously in this system(a domain wall between the two vacua), we arrive at a new phase whose category of symmetric defects is exactly $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$. This because when time-reversal is broken, the $T\in{\mathbb{Z}}_{2}^{T}$-action creates a nontrivial IR defect in the system that fuses with defects in $\mathcal{C}$ according to the ${\mathbb{Z}}_{2}^{T}$-action on $\mathcal{C}$. The process of going from an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$ to the Morita equivalent Galois-real fusion category $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$ describes the spontaneously breaking of time-reversal.
@@ V.1 The Setup
The action of $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$ on the category of boundary conditions $\mathcal{M}$ provides $\mathcal{M}$ with a module category structure: $\mathcal{C}\times\mathcal{M}\to\mathcal{M}$ that is only $\mathbb{R}$-linear: If $a$ is an anti-linear line, then a scalar on the boundary gets complex conjugated when it passes the junction. This means the functor
@@
For any symmetry category $\mathcal{C}$, one can define a 2-category as follows.
  • Objects are (1+1)(1+1)d gapped phases enriched with $\mathcal{C}$.
@@
The 2-category of (1+1)(1+1)d gapped phases enriched with $\mathcal{C}$ is $\mathsf{Mod}(\mathcal{C})$.
For instance, the 2-category of (1+1)(1+1)d gapped phases with ${\mathbb{Z}}_{2}^{T}$ symmetry is $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})\simeq 2\mathsf{Vec}_{\mathbb{R}}$ which has three simple objects: $\mathsf{Vec}_{\mathbb{R}},\mathsf{Vec}_{\mathbb{C}},\mathsf{Vec}_{\mathbb{H}}$.
@@ V.2 Order Parameters and Drinfeld Center
As we have discussed for time-reversal in Section II.3.2, a pure time-reversal symmetry does not have a string order that can detect the Haldane phase. This is not special to time-reversal: any anti-linear symmetry generator does not have a string order parameter. We will argue that given this observation, the Drinfeld center $\mathcal{Z}(\mathcal{C}_{1}\oplus\mathcal{C}_{T})$ is still the correct category of all local and string order parameters of an anti-linear symmetry $\mathcal{C}_{1}\oplus\mathcal{C}_{T}$. Drinfeld center is a notion that can be defined for any monoidal category, hence well-defined for real fusion categories, and in particular for Galois-real fusion categories. For a categorical anti-linear symmetry $\mathcal{C}:=\mathcal{C}_{1}\oplus\mathcal{C}_{T}$, a simple object in $\mathcal{Z}(\mathcal{C})$ is again given by a pair $(x,\gamma)$, where $x\in\mathcal{C}$ and $\gamma:x\otimes y\simeq y\otimes x,\penalty 10000\ \forall y\in\mathcal{C}$ is the half-braid. However, as shown in [15], the object $x$ is in fact strictly in $\mathcal{C}_{1}$, otherwise there is no consistent half-braid $\gamma$. This is because, assuming a half-braid $\gamma$ does exists, then it gives natural isomorphism $x\otimes\mathbf{1}\simeq\mathbf{1}\otimes x$, which identifies the left and right actions by scalars, implying $x$ must be a linear(Galois-trivial) object $x\in\mathcal{C}_{1}$. This precisely matches with the expectation that an anti-linear symmetry generator does not have a string order parameter: assume a string order parameter does exist, then it can be used to identify the action by scalars from two sides of the string order:
@@
Here we are not assuming any SymTFT interpretation of the Drinfeld center, simply viewing it as the category of 1D order parameters associated with $\mathcal{C}$. The SymTFT interpretation will be discussed in Section VI. However it is already clear from the discussion here that the Drinfeld center can not describe gapped phases that are not captured by string orders, such as the Haldane chain.
@@ V.3.1 Gauging between Symmetry Categories
Recall that for two complex fusion categories $\mathcal{C},\mathcal{D}$, we say they are Morita equivalent if there is some algebra $A\in\mathcal{C}$ such that $\mathcal{D}\simeq\mathsf{Bimod}_{\mathcal{C}}(A)$ are equivalent as fusion categories. In this case we say $\mathcal{D}$ is obtained by condensing $A$ in $\mathcal{C}$. Equivalently, $\mathcal{C},\mathcal{D}$ are Morita equivalent if there is a $\mathcal{C}$-module category $\mathcal{M}$ such that $\mathcal{D}\simeq\text{End}_{\mathcal{C}}(\mathcal{M})$. If $\mathcal{C},\mathcal{D}$ are symmetry categories for two internal generalized symmetries, then being Morita equivalent has the consequence that the 2-categories of gapped phases with symmetry $\mathcal{C}$ and $\mathcal{D}$: $\mathsf{Mod}(\mathcal{C}),\penalty 10000\ \mathsf{Mod}(\mathcal{D})$ are equivalent. This means we can relabel phases enriched with symmetry $\mathcal{C}$ as phases enriched with symmetry $\mathcal{D}$ in such a way that does not affect any physical observables.
The Morita theory works well over $\mathbb{R}$ as well, see our discussion in Section III.3. Hence we say two anti-linear symmetry categories $\mathcal{C},\mathcal{D}$ are dual to each other if they are Morita equivalent as real fusion categories. This means there exists some algebra $A\in\mathcal{C}$ such that $\mathcal{D}\simeq\mathsf{Bimod}_{\mathcal{C}}(A)$. Notice that since both $\mathcal{C},\mathcal{D}$ are Galois-real, the algebra $A$ is necessarily in $\mathcal{C}_{1}$ according to Theorem III.4. This is the mathematical manifestation of the fact that one cannot literally gauge time-reversal, or any anti-linear symmetry generator. The gauging always happens inside the linear part $\mathcal{C}_{1}$. Then if $\mathcal{C},\mathcal{D}$ are two dual anti-linear symmetries, the 2-category of gapped phases: $\mathsf{Mod}(\mathcal{C}),\penalty 10000\ \mathsf{Mod}(\mathcal{D})$ are equivalent.
@@ V.4 ${\mathbb{Z}}_{2}^{T}$ Gapped Phases Revisited
We discussed the three ${\mathbb{Z}}_{2}^{T}$ gapped phases in Section II including their module category descriptions. One thing that was not discussed there is the underlying gapped phase. Let us confirm here that the three phases have the expected underlying gapped phase.
@@ V.5 General Structure of $G^{T}$ Gapped Phases
We now consider (1+1)(1+1)d gapped phases with a general group-like $G^{T}$-symmetry and with no anomaly. The symmetry category is $\mathsf{Vec}_{G^{T}}$ and $G^{T}$-gapped phases are module categories over $\mathsf{Vec}_{G^{T}}$. We have classified modules over $\mathsf{Vec}_{G^{T}}$ in Section IV.1.2. The indecomposable modules are denoted as $\mathcal{M}(K^{T},\omega)$ where $K^{T}<G^{T}$ and $\omega$ is a twisted 2-cocycle for $K^{T}$. It is clear that this matches with the standard physical classification of (1+1)(1+1)d gapped phases with $G^{T}$-symmetry. The subgroup $K^{T}$ stands for the unbroken symmetry in the phase, and $\omega$ labels the $K^{T}$-SPT phase in any of the vacua. From the general discussion in Section V.1, the number of ground states in a phase defined by $\mathcal{M}(K^{T},\omega)$ is the number of simple objects in $\mathcal{M}(K^{T},\omega)$, which is $|G^{T}/K^{T}|$.
According to the general discussion in Section V.1, the category of symmetric defects in the phase $(K^{T},\omega)$ is given by the Morita dual real fusion category
@@
Then by Theorem III.4, this real fusion category is Galois-real exactly when $K^{T}\leq G_{0}$ is completely unitary and $\mathbb{R}$-real when $K^{T}$ contains any anti-unitary element. This also means the Morita dual categories $\mathsf{Bimod}_{\mathsf{Vec}_{G^{T}}}({\mathbb{C}}[K^{T}]^{\omega})$ for $K^{T}\leq G_{0}$ can be viewed as symmetries dual to $\mathsf{Vec}_{G^{T}}$.
@@
In the remainder of this section we study several concrete examples of gapped phases with anti-linear symmetries.
@@ V.6 ${\mathbb{Z}}_{4}^{T}$-phases
The symmetry categories $\mathsf{Vec}_{{\mathbb{Z}}_{4}^{T}}$, $\mathsf{Vec}_{{\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}^{T}}^{\alpha}$ are Morita equivalent.
Here $\alpha$ is a nontrivial 3-cocycle given in Equation V.15. We notice this is similar to the situation with an internal ${\mathbb{Z}}_{4}$-symmetry, where it is Morita equivalent to an anomalous ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$-symmetry.